OpenStudy (shargrett):

Which of the following is a polynomial with roots: −square root of 3, square root of 3, and 2? asap help!!

OpenStudy (sshayer):

if the roots are a,b&c then the polynomial in x is (x-a)(x-b)(x-c) \[\left( x-\sqrt{a} \right)\left( x+\sqrt{a} \right)=x^2-\left( \sqrt{a} \right)^2=x^2-a\]